# pnasresearcharticle

### v1

@@ Discussion
Our individual mobility habits are critical to our transport system implicating congestion, environmental pollution, and the travel experience. There is currently no empirical measure that can capture individual mobility habits at scale. The current study implements the first such measure capturing the habitual aspects of day-to-day departure time distribution. Measuring the time distribution of public transit riders’ tendency towards choosing the same departure time repeatedly—that we term ‘stickiness’ (27) is important in its capacity to provide the first empirical measure of travel habits of individual users across a metropolis.
Our measure of departure time peakedness comprises a set of three related measures that collectively capture a user’s tendency towards repeatedly choosing [or ‘sticking’] the same departure time for a given origin and destination within the bus network. High departure time peakedness indicates users ‘sticking’ to a preferred departure time, with low peakedness capturing users with less preference towards a particular time of departure.
For public transit operators, considering individual users’ departure time peakedness is important in understanding the relative impact of a service changes/disruptions. Given that users associated with high departure time peakedness are typically associated with higher rescheduling penalties and lower willingness to shift departure time, service changes/disruptions creating departure time changes will likely to impact users with stronger departure time peakedness than on those with lower peakedness levels.
@@
For metropolitan councils and policy makers, understanding how broader changes in land use, transport policy and pandemics act to re-shape users’ departure time peakedness are important in helping to inform future initiatives – such as the likely implication on users in removal, modification or introduction of new routes and services. Particular stops and routes associated with higher levels of peakedness are likely to induce more disruption if shifts to service scheduling and routes were imposed compared to stops and routes with users characterised by lower peakedness levels.
@@
[…]
The authors would like to thank Translink for the supply of the smart card data on which the current study is based.
[…]
[2]
Supporting Information (SI)
1 Methods
In this study, we define the ‘peakedness’ of a departure time distribution as the extent to which departure time observations are concentrated around the peak period or the most preferred departure time. When applied to a user’s day-to-day departure time choices, departure time peakedness quantifies the user’s tendency towards choosing the same departure time repeatedly for a given trip. A user with a strong preference for a particular departure time will exhibit a departure time distribution with a high degree of peakedness, characterised by a sharp concentration around the user’s preferred departure time. Conversely, a user with more variable departure time choices will have a flatter, less peaked distribution, reflecting greater departure time flexibility.
1.A Limitations of Existing Measures
@@
Kurtosis is another common measure that describes the shape of a distribution, particularly its ‘tailedness’ and the sharpness of its peak. While it provides some insight into whether a distribution has a sharp or flat peak, the interpretation of kurtosis in terms of peakedness is incorrect because it does not directly measure the shape of the peak (‘peakedness’)—the regularity or consistency of repeated observations around the peak)—but primarily describes the shape of the tails (‘tailedness’) (32). Moreover, kurtosis is sensitive to outliers and extreme values in the tails of the distribution, which can inflate or distort the measure even if the central concentration of data remains strong.
@@
1.B Birnbaum’s Concept of Relative Peakedness
In his 1948 paper ‘On Random Variables with Comparable Peakedness’ (33), Birnbaum introduced the concept of ‘relative peakedness’ to correctly measure what is intuitively considered as the peakedness of a distribution and evaluate the concentration of probability distributions around their peaks. The key idea behind relative peakedness is to quantify how sharply or strongly a distribution is concentrated around its mode or peak ’relative to’ other distributions. That is, it assesses how much one distribution is more ‘peaked’ compared to another reference distribution, rather than evaluating the peakedness of a distribution in isolation. Birnbaum’s definition of relative peakedness is presented below:
Definition 1 (Relative Peakedness (33)).
@@ 1.B Birnbaum’s Concept of Relative Peakedness
1.C Proposed Peakedness Metric for Departure Time Distribution
A limitation of Birnbaum’s relative peakedness is that it only establishes a comparative condition between two distributions and does not provide a direct measure to quantify the peakedness of an individual distribution. To address this limitation, we extend Birnbaum’s concept to develop a quantitative measure that characterises the peakedness of a single probability distribution. Inspired by the idea of assessing the peakedness of a distribution based on ‘the concentration of data points within a specified range’, we measure the degree to which data points are concentrated within the most peaked range of an individual distribution, focusing on ‘the highest concentration of data points’, rather than relative concentration of data points as in Birnbaum’s approach. By applying this concept to travellers’ departure time distributions, we propose the $h$-minute Peak Trip Concentration (PTC) metric, denoted by $\psi_{h}$, which measures how much departure times are concentrated within the most peaked $h$-minute window of the day. This provides a direct measure to quantify the peakedness of an individual departure time distribution. The definition of $\psi_{h}$ is presented below:
Definition 2 ($h$-minute Peak Trip Concentration).
Let $X$ be a real random variable representing the departure time for a given trip, and let $t$ be a real constant representing a specific time of day, where both $X$ and $t$ are expressed in total minutes since the start of the day. For a given time window of $h$ minutes, we define the $h$-minute peak trip proportion, denoted by $\psi_{h}$, as the maximum probability that the departure time $X$ falls within any $h$-minute window. Mathematically, this is expressed as:
[…]
where $f_{X}(x)$ is the probability density function of $X$.
Eq.[8] finds the time window $[t,t+h]$ that contains the largest proportion of departure times, and the corresponding interval probability $\psi_{h}$ represents the proportion of trips that fall within the most peaked $h$-minute window. Although the proposed metric fundamentally measures the probability of events occurring within a defined window—something not new in probability theory—, the novelty lies in its application as a peakedness metric for departure time distributions. As mentioned in Section 1.A, traditional measures such as standard deviation, kurtosis, or entropy assess the overall characteristics of a distribution but may not fully capture how sharply or consistently departure times are concentrated within a specific interval. The proposed metric addresses this gap by focusing on the maximum concentration within an $h$-minute window, directly measuring the peakedness of the departure time pattern. By quantifying the probability that departures cluster around a specific peak time, the proposed metric offers a robust measure of behaviour that is less sensitive to outliers. This is a significant advantage over traditional statistical measures, which consider the entire range of the distribution. Therefore, while the concept of measuring probabilities is familiar, applying it to assess individual peakedness in departure time distributions represents a novel and practical contribution.
1.D Conceptual Similarities with Peak Hour Factor (PHF)
@@
1.E Choice of Peak Window ($h$)
Selecting the appropriate width of the peak window, $h$, is important in measuring and interpreting $\psi_{h}$, particularly when comparing its values across travellers. If $h$ is too large, $\psi_{h}$ will tend to approach 1, as most departures fall within the wide window, leading to an inflated peakedness for most travellers. Conversely, if $h$ is too small, $\psi_{h}$ will approach 0, since very few departures times fall within the narrow window, resulting in artificially low peakedness across users. In both cases, using an overly large or small $h$ diminishes the ability to differentiate between varying levels of departure time peakedness. Manually selecting an appropriate $h$ can be difficult, so an automatic method based on data is preferable. The ideal $h$ should allow $\psi_{h}$ to effectively distinguishes between different users’ departure time patterns. One way to achieve this is by analysing the variance of $\psi_{h}$ across users. When $h$ is very small, the variance will be low because $\psi_{h}\approx 0$ for most users. As $h$ increases, the variance initially rises but eventually decreases again when $h$ becomes too large and $\psi_{h}\approx 1$ for most users. Therefore, the variance of $\psi_{h}$ as a function of $h$ forms a peak at a critical point, $h=h^{*}$, after which it decreases as depicted in Figure 9(a). This critical point, $h=h^{*}$, serves as the optimal window width that maximises the spread of $\psi_{h}$ values across users, ensuring the greatest distinction between different departure time peakedness levels. Given a dataset of departure time distributions from $n$ users, the optimal window width $h^{*}$ can be determined by finding the value of $h$ that maximises the variance of $\psi_{h}$ across the $n$ users, as follows:
where $\psi_{h,i}$ is the $h$-min peak trip concentration of user $i$, $\overline{\psi_{h}}$ is the sample mean of $\psi_{h}$ over $n$ users, and $\text{Var}[\psi_{h}]$ is the variance of $\psi_{h}$, which can be estimated ($\mathrel{\hat{=}}$) by calculating the sample variance. The $h$-min peak trip concentration under this optimal $h^{*}$, $\psi_{h^{*}}$, can be seen as the most discriminative, as it maximises the spread of $\psi_{h}$ values across users, thereby best distinguishing different levels of peakedness in the system.
Figure 7 shows the estimation results of $\psi_{h}$ for different $h$ values, namely $h=5,10,\cdots,40$ minutes, for Brisbane and Gold Coast during two time-of-day periods (AM and PM). For each $h$, the distribution of the measured $\psi_{h}$ values across users is shown as a histogram, along with its mean ($\overline{\psi_{h}}$) and variance (Var). As expected, the distribution of the measured $\psi_{h}$ values shift to the right (i.e., the mean, $\overline{\psi_{h}}$, increases) as the window width, $h$, becomes larger. Meanwhile, the diversity or spread of the measured $\psi_{h}$ values (i.e., $\text{Var}[\psi_{h}]$) initially increases and then decreases as $h$ increases. Figure 8 provides a separate plot showing the change in $\text{Var}[\psi_{h}]$ with respect to $h$, where the $h$ value that maximises the variance (optimal peak window, $h^{*}$) is marked by a vertical dashed line.
The analysis reveals that the value of $h^{*}$ is consistent across the three study networks. Specifically, all three network indicate that $h^{*}$ is 15 minutes for the AM period and $h^{*}$ is 20-25 minutes for the PM period. The shorter optimal window for the AM period likely reflects more concentrated departure time distributions, as morning schedules are typically more constrained due to fixed start times for work or school. In contrast, departure times may be more spread out due to the varying end times of activities. When many individuals have tightly concentrated departure times (as in the AM period), the variance of $\psi_{h}$ values across users will peak at a shorter $h$ because smaller windows are sufficient to capture the high concentration of departures for most users. On the other hand, when the population has more dispersed or spread-out departure times (as in the PM period), the variance of $\psi_{h}$ across the population will reach its peak at a larger $h^{*}$ value because a larger $h$ window is needed to capture significant portions of the departure events across users.
Despite the differences in the optimal window values between the AM and PM periods, the variance around those values is not significantly different. Given this, and for the sake of consistent comparison and interpretation of peakedness measures ($\overline{\psi_{h}}$, $\psi_{h}^{sys}$, and $PCF_{h}$) across different cases, we propose selecting a fixed window of $h=20$ minutes for this study. This approach allows for uniform analysis while still adequately capturing the key characteristics of both AM and PM departure time distributions.
Figure 7: Histograms of estimated $h$-min peak trip concentration ($\psi_{h}$) across users under different $h$ values ($h=5,10,\cdots,40$ minutes): the mean ($\overline{\psi_{h}}$) and variance (Var) of $\psi_{h}$ are presented in each histogram.
Figure 8: Distribution of estimated $h$-min peak concentration ($\psi_{h}$) across users under different $h$ values ($h=5,10,\cdots,40$ minutes): $h$ that produces the most dispersed $\psi_{h}$ distribution (i.e., maximising $Var[\psi_{h}]$) is shown in red box ( $h^{*}$ ).
1.F Further Analysis of Optimal Peak Window ($h^{*}$)
@@
To get some intuition about what the variance-maximising $h^{*}$ in Eq.(9) represents, we derive $h^{*}$ in a simplified system consisting of two types of users—user group 1 with high peakedness and user group 2 with low peakedness. Figure 9(b) illustrates the probability density curves for these two user groups, where $f_{1}$ and $f_{2}$ represent the departure time distribution of high peakedness users and low peakedness users, respectively. The $h$-min peak concentration under $f_{1}$ and $f_{2}$, denoted by $\psi_{h,1}$ and $\psi_{h,2}$, are also shown as the shaded areas under the curves. In this setting, we have found that $h^{*}$ corresponds to the peak window at which the two density curves intersect as indicated by the two vertical dotted lines in Figure 9(b).
Figure 9: Illustration of optimal h.
@@
To prove this, we show that the variance of $\psi_{h}$ increases, i.e., $\text{Var}[\psi_{h}]<\text{Var}[\psi_{h^{\prime}}]$ for $h<h^{\prime}$, when $h$ and $h^{\prime}$ are less than or equal to $h^{*}$, where we show $\text{Var}[\psi_{h}]<\text{Var}[\psi_{h^{\prime}}]$ occurs only under the condition $\psi_{h^{\prime},1}-\psi_{h,1}>\psi_{h^{\prime},2}-\psi_{h,2}=\Delta\psi_{h,1}>\Delta\psi_{h,2}$ that happens when the $f_{1}$ curve (red) is above the $f_{2}$ curve (blue) as illustrated in Figure 9(c) ($h,h^{\prime}\leq h^{*}$). Similarly, we show that the variance decreases, i.e., $\text{Var}[\psi_{h}]\geq\text{Var}[\psi_{h^{\prime}}]$ for $h<h^{\prime}$, when $h$ and $h^{\prime}$ are greater than or equal to $h^{*}$, where $\text{Var}[\psi_{h}]\geq\text{Var}[\psi_{h^{\prime}}]$ occurs only under the condition $\psi_{h^{\prime},1}-\psi_{h,1}\leq\psi_{h^{\prime},2}-\psi_{h,2}=\Delta\psi_{h,1}\leq\Delta\psi_{h,2}$ that happens when the $f_{1}$ curve (red) is below the $f_{2}$ curve (blue) as illustrated in Figure 9(d) ($h,h^{\prime}\geq h^{*}$). With these, we can see that the variance of $\psi_{h}$ is maximised at $h=h^{*}$, where $h^{*}$ is the boundary between the two conditions, $\Delta\psi_{h,1}>\Delta\psi_{h,2}$ and $\Delta\psi_{h,1}\leq\Delta\psi_{h,2}$. This property is formally described by Proposition 1 below, followed by the detailed proof.
@@
2 Data
The three peakedness measures developed in this work—$\psi_{h}$, $\psi_{h}^{sys}$, and $PCF_{h}$—are estimated using public transport smart card data from South East Queensland (SEQ), Australia, to characterise the departure time stickiness of bus passengers. This study considers three different study networks, namely, Brisbane, Gold Coast, and Sunshine Coast, which are three major cities in the SEQ region.
We use a 12-month (July 2015 to June 2016) record of go card, an electronic smart card ticketing system used on the public transport network in SEQ operated by TransLink (the public transit agency for Queensland). The go card is the principal means by which passengers pay their transit fares in SEQ, representing approximately 87% of all trips taken across the TransLink public transport network (34). It requires passengers to touch the card on a card reader at the start and end of each journey as well as when transferring between trips, where TransLink defines a trip as the act of travelling from point A to point B with no transfers (a single trip) and a journey as the act of travelling from the origin to the final destination that may involve one or a number of trips (transfers) using different transport modes (35). The go card data contain information on each trip’s boarding and alighting locations and their associated times, trip ID, journey ID, card ID (go card identifier), and passenger type, among other attributes. Passenger types recorded in the go card database include Adult, Senior/Pensioner, Tertiary Student, Secondary Student (secondary school students aged 15+ years), Child (children aged 5 to 14 years), and others (special concession card holders such as veterans).
In this study, we focus on regular bus passengers who use buses repeatedly. For each bus user, trip data are processed and grouped to form journeys for each origin-destination (OD). Then, the departure time distribution is constructed for each user-OD pair by extracting the boarding times at the origin stop for a given OD journey over the course of one year from July 2015 to June 2016. The user-OD pair is, thus, the basic unit for measuring the peakedness of user-level departure time distribution, i.e., user-level $h$-min peak proportion ($\psi_{h}$), as shown in the examples in Figure 10.
Figure 10 shows examples of three bus users’ departure time distributions and the estimated $\psi_{h}$ for a peak window width of $h=15$ minutes. The $\psi_{h}$ values for Users 1, 2, and 3 are 0.843, 0.445, and 0.152, respectively. User 1 has the strongest tendency towards choosing a particular departure time as the great majority of his/her trips (84.3%) occur during the short 15-min peak window. User 2 shows more flexibility in choosing departure time, resulting in about half his/her trips (44.5%) occurring during the 15-min peak. User 3 has the lowest stickiness and the greatest flexibility among three, showing that his/her most peaked 15-minutes only covers 15.2% of the trips and the remaining 84.4% of data are widely spread out from this peak.
Table 2 shows the summary of dataset size for the three study networks after filtering in terms of the number of users, user-OD pairs, journeys, and origins, with and without grouping by passenger type. The number of users is the total count of all distinct go card users. Since each user may have multiple ODs and, thus, result in multiple user-OD pairs, the number of distinct user-OD pairs is greater than the number of users. The number of journeys is the sum of the number of journeys across all user-OD pairs. The number of origins is the number of unique bus stops from which these journeys started. Overall, the Brisbane dataset is the largest, which has 50,336 user-OD pairs providing 50,336 departure time distributions and their associated ($\psi_{h}$) measures, followed by Gold Coast and Sunshine Coast, which has 3968 and 1817 user-OD pairs, respectively. In terms of passenger type, Adult is the largest group in all three cities. The second largest group differs, however, as Tertiary Student is the second largest group in Brisbane and Gold Coast, whereas Senior/Pensioner is the second largest group in Sunshine Coast.
Table 2: Summary of datasets
3 Results
Trend analysis
An interesting question in characterising departure time peakedness is whether system- or user-level peakedness characteristics evolve over time. To explore this, we measured both $\psi_{h}^{sys}$ and $\psi_{h}$, with $h$ = 20 minutes, using data from a three-month moving window spanning over one year from July 2015 to June 2016. This results in a time-series with 10 data points, corresponding to the periods July-September 2015, August-October 2015, and so on, up to April-June 2016. These data points allow us to analyse trends or changes in the chosen peakedness measure over time.
We then applied the Mann-Kendall (MK) Trend Test (30, 31) to detect significant trends in time series. As our time series data are relatively short, we used the original non-parametric MK test for analysing consistently increasing or decreasing trends (monotonic trends), which does not consider serial correlation or seasonality effects (36). The MK test determines whether to reject the null hypothesis of no monotonic trend at a given significance level $\alpha$, which is set to 0.05 in this study.
For the system-level peakedness, the MK test results show that there is no monotonic trend in the series, which is true for all three cities for both AM and PM. Figure 11 shows the time-series of $\psi_{h}^{sys}$ used in the MK test, where the x-axis labels the first month of each three-month moving window in the format ‘YY/MM’ (Year/Month). The number of long-term regular users involved in constructing each system-level time-series ($n$) is shown in the parenthesis next to each city name within the legend. In all six cases, the value of $\psi_{h}^{sys}$ remains quite stable over the one-year period, indicating the absence of a significant trend in the system-level departure time peakedness. In contrast, the MK test results reveal interesting trends in the user-level peakedness time-series data. Approximately 25% of the long-term regular users exhibit statistically significant trends in their $\psi_{h}$ time-series. Of these users, half show increasing trends and the other half show decreasing trends. This pattern is consistent across all three cities, regardless of the selected peak window width ($h$). The detailed results are displayed in Figure 12, which shows the proportion of user-OD pairs with decreasing or increasing $\psi_{h}$, both as a percentage of long-term regular users (indicated by taller bars in the background) and as a percentage of the original regular user base (indicated by shorter bars in the foreground). The results are obtained for the three cities using four different peak window widths ($h$ = 5, 10, 20, and 30 minutes). The first sub-figure shows the number of user-OD pairs exhibiting any monotonic trends (either decreasing or increasing $\psi_{h}$), while the second and third sub-figures show the numbers corresponding to increasing and decreasing trends, respectively.

### v2

@@ 4 Conclusion
Our individual mobility habits are critical to our transport system implicating congestion, environmental pollution, and the travel experience. There is currently no empirical measure that can capture individual mobility habits at scale. The current study implements the first such measure capturing the habitual aspects of day-to-day departure time distribution. Measuring the time distribution of public transit riders’ tendency towards choosing the same departure time repeatedly is important in its capacity to provide the empirical measure of travel habits of individual users across a metropolis. Our measure of departure time peakedness comprises a set of three related measures that collectively capture a user’s tendency towards repeatedly choosing the same departure time for a given origin and destination within the bus network. High departure time peakedness indicates users ‘sticking’ to a preferred departure time, with low peakedness capturing users with less preference towards a particular time of departure.
Our findings demonstrate that departure time peakedness is for the main part driven by intrinsic passengers characteristics (captured in the current study by passenger type: adults, students, children, seniors and pensioners), rather than external variables such as weather and holiday periods. Furthermore, while individual passengers show significant variation in their departure time peakedness over extended periods of time (offering some evidence to suggest that personal travel habits may transform gradually), at the overall system-level, peakedness exhibits remarkably long-term stability. Another key insight from our analysis is that system-level departure time peakedness does not always reflect individual user-level patterns. Traditional transit planning and operations often rely on system-level departure time distributions to identify peak hours and assess demand patterns. However, this information alone does not offer insights into individual travel behaviours, such as users’ responses towards changes or disruptions in public transport services. Distinguishing user-level peakedness from system-level peakedness is crucial for assessing disruption impacts and developing targeted mitigation strategies. For public transit operators, considering individual users’ departure time peakedness is important in understanding the relative impact of a service changes/disruptions. Given that users associated with high departure time peakedness are typically associated with higher rescheduling penalties and lower willingness to shift departure time, service changes/disruptions creating departure time changes will likely to impact users with stronger departure time peakedness than on those with lower peakedness levels.
From a managerial perspective, the choice of the peak window width $h$ (in the current study that was optimally defined at 20 minutes) has important implications. For example, a smaller $h$ (e.g., 5 minutes) allows for the identification of extremely rigid travellers. Since the proportion of trips occurring within such a narrow window is generally small, users with high $h$-min peak concentrations within this small window indicate a very rigid and consistent departure time preference. Conversely, a larger $h$ (e.g., 30 minutes) allows us to identify particularly flexible travellers. Since trip proportions within a larger window are generally higher, users with low $h$-min peak concentration in this context can be considered as having particularly more dispersed departure time patterns, potentially less vulnerable to service disruptions. For strategic planning, a smaller $h$ might help focus on the most intense peaks, while a larger $h$ would have value in informing boarder capacity planning decisions. In demand management, exploring changes in peak trip concentration before and after a major intervention—such as the introduction of flexible working arrangements by a major employer, the commencement of congestion charging or the introduction of highly subsided fares (as seen in SEQ)—can reveal the extent to which policies resulted in increasing or redistributing demand or simply shifting the peak window. From an operational perspective, the choice of the peak window width could be aligned to operational capabilities, such that for urban transit, the 20-minute window (that we employed in this study) might be most appropriate, whereas for other part of the system (such as the airport transit services) or for weekend operations a larger window width may work better.
@@
For metropolitan councils and policy makers, understanding how broader changes in land use, transport policy and pandemics act to re-shape users’ departure time peakedness are important in helping to inform future initiatives. This might include the likely implication on users following the removal, modification or introduction of new routes and services. Particular stops and routes associated with higher levels of peakedness are likely to induce more disruption if shifts to service scheduling and routes were imposed compared to stops and routes with users characterised by lower peakedness levels.
@@
Our framework can be extended in various ways to further explore habitual travel behaviour and its implications. First, while we focus on smart card data for bus passengers, the peakedness measurement can be applied to any departure or arrival time data, such as Global Positioning System (GPS) records, to analyse broader traveller populations, including private cars and freight. This could reveal different departure or arrival time peakedness patterns across modes with varying levels of flexibility and constraints. Second, integrating richer activity, demographic, and socio-economic data—such as trip purpose, income, occupation, or household structure—could provide deeper insights into how trip context and personal circumstances influence habitual travel behaviours. Finally, studying how departure time peakedness changes in response to major disruptions (e.g., pandemics, extreme weather, or infrastructure changes) could enhance our understanding of travel behaviour resilience and adaptation. A critical implication is that enabling such investigations requires longitudinal day-to-day observations of repeated trip choices over an extended period to capture proper choice distributions. Single-day travel data, like traditional household travel surveys, are insufficient for in-depth analysis of habitual behaviours. Expanding access to longitudinal mobility data through smart card systems, GPS, Internet-of-Things (IoT) devices, and other emerging technologies is essential for advancing research in this area.
@@
Appendix A Limitations of Existing Measures
@@
Kurtosis is another common measure that describes the shape of a distribution, particularly its ‘tailedness’ and the sharpness of its peak. While it provides some insight into whether a distribution has a sharp or flat peak, the interpretation of kurtosis in terms of peakedness is incorrect because it does not directly measure the shape of the peak (‘peakedness’)—the regularity or consistency of repeated observations around the peak)—but primarily describes the shape of the tails (‘tailedness’) (Westfall, 2014). Moreover, kurtosis is sensitive to outliers and extreme values in the tails of the distribution, which can inflate or distort the measure even if the central concentration of data remains strong.
@@
Appendix B Birnbaum’s Concept of Relative Peakedness
In his 1948 paper ‘On Random Variables with Comparable Peakedness’ Birnbaum (1948), Birnbaum introduced the concept of ‘relative peakedness’ to correctly measure what is intuitively considered as the peakedness of a distribution and evaluate the concentration of probability distributions around their peaks. The key idea behind relative peakedness is to quantify how sharply or strongly a distribution is concentrated around its mode or peak ’relative to’ other distributions. That is, it assesses how much one distribution is more ‘peaked’ compared to another reference distribution, rather than evaluating the peakedness of a distribution in isolation. Birnbaum’s definition of relative peakedness is presented below:
Definition 1 (Relative Peakedness Birnbaum (1948)).
@@ Appendix B Birnbaum’s Concept of Relative Peakedness
A limitation of Birnbaum’s relative peakedness is that it only establishes a comparative condition between two distributions and does not provide a direct measure to quantify the peakedness of an individual distribution. To address this limitation, we extend Birnbaum’s concept to develop a quantitative measure that characterises the peakedness of a single probability distribution. Inspired by the idea of assessing the peakedness of a distribution based on ‘the concentration of data points within a specified range’, we measure the degree to which data points are concentrated within the most peaked range of an individual distribution, focusing on ‘the highest concentration of data points’, rather than relative concentration of data points as in Birnbaum’s approach. By applying this concept to travellers’ departure time distributions, we propose the $h$-minute Peak Trip Concentration (PTC) metric, $\psi_{h}$ in Eq. [1], which measures how much departure times are concentrated within the most peaked $h$-minute window of the day. This provides a direct measure to quantify the peakedness of an individual departure time distribution.
Appendix C Conceptual Similarities with Peak Hour Factor
@@
Appendix D Further Analysis of Optimal Peak Window ($h^{*}$)
@@
To get some intuition about what the variance-maximising $h^{*}$ in Eq.(7) represents, we derive $h^{*}$ in a simplified system consisting of two types of users—user group 1 with high peakedness and user group 2 with low peakedness. Figure 17(a) illustrates the probability density curves for these two user groups, where $f_{1}$ and $f_{2}$ represent the departure time distribution of high peakedness users and low peakedness users, respectively. The $h$-min peak concentration under $f_{1}$ and $f_{2}$, denoted by $\psi_{h,1}$ and $\psi_{h,2}$, are also shown as the shaded areas under the curves. In this setting, we have found that $h^{*}$ corresponds to the peak window at which the two density curves intersect as indicated by the two vertical dotted lines in Figure 17(a).
Figure 17: Illustration of optimal peak window: the interpretation of variance-maximising $h^{*}$ in a system with two departure time peakedness levels
@@
To prove this, we show that the variance of $\psi_{h}$ increases, i.e., $\text{Var}[\psi_{h}]<\text{Var}[\psi_{h^{\prime}}]$ for $h<h^{\prime}$, when $h$ and $h^{\prime}$ are less than or equal to $h^{*}$, where we show $\text{Var}[\psi_{h}]<\text{Var}[\psi_{h^{\prime}}]$ occurs only under the condition $\psi_{h^{\prime},1}-\psi_{h,1}>\psi_{h^{\prime},2}-\psi_{h,2}=\Delta\psi_{h,1}>\Delta\psi_{h,2}$ that happens when the $f_{1}$ curve (red) is above the $f_{2}$ curve (blue) as illustrated in Figure 17(b) ($h,h^{\prime}\leq h^{*}$). Similarly, we show that the variance decreases, i.e., $\text{Var}[\psi_{h}]\geq\text{Var}[\psi_{h^{\prime}}]$ for $h<h^{\prime}$, when $h$ and $h^{\prime}$ are greater than or equal to $h^{*}$, where $\text{Var}[\psi_{h}]\geq\text{Var}[\psi_{h^{\prime}}]$ occurs only under the condition $\psi_{h^{\prime},1}-\psi_{h,1}\leq\psi_{h^{\prime},2}-\psi_{h,2}=\Delta\psi_{h,1}\leq\Delta\psi_{h,2}$ that happens when the $f_{1}$ curve (red) is below the $f_{2}$ curve (blue) as illustrated in Figure 17(c) ($h,h^{\prime}\geq h^{*}$). With these, we can see that the variance of $\psi_{h}$ is maximised at $h=h^{*}$, where $h^{*}$ is the boundary between the two conditions, $\Delta\psi_{h,1}>\Delta\psi_{h,2}$ and $\Delta\psi_{h,1}\leq\Delta\psi_{h,2}$. This property is formally described by Proposition 1 below, followed by the detailed proof.
@@
Appendix E Departure Time Peakedness vs. Arrival Time Peakedness
An analysis was conducted to compare how the peakedness measure of a user may differ between departure times at the origin and arrival times at the destination. For each individual user in the Brisbane dataset, the $h$-minute Peak Trip Concentration, $\psi_{h}$, was estimated based on both the departure time distribution and the arrival time distribution for a given origin-destination (OD) trip using $h$=20 minutes. Figure 18 shows a scatter plot where each blue dot represents a pair of $\psi_{h}$ values—one from the departure time distribution and the other from the arrival time distribution for the same user. The results indicate a strong positive correlation between departure time peakedness and arrival time peakedness, as evidenced by the nearly 45-degree fitted linear regression line (shown in red): $y=0.91+0.02$ ($R^{2}=0.94$). This finding suggests that, despite additional factors affecting arrival times such as congestion, travel time variations, and route changes, the degree of peakedness in arrival times remains largely consistent with that of departure times. In other words, users who exhibit high departure time peakedness also tend to have have high arrival time peakedness. This supports the validity and robustness of using departure time peakedness as a meaningful indicator of individual travel routines to study habitual behaviours, without requiring to explicitly factor in arrival distributions.
Figure 18: A scatter plot of $h$-minute Peak Trip Concentration, $\psi_{h}$, for departure time vs. arrival time distributions for individual users in Brisbane ($h$=20 minutes). The strong correlation between departure time peakedness and arrival time peakedness is demonstrated by the nearly 45-degree linear fit line (red).
Appendix F Destination Zones with High Arrival Time Peakedness
In addition to analysing origin zones with high departure time peakedness, as shown in Figure 16 in Section 3.6, we conducted a complementary analysis of destination zones by estimating the three peakedness measures $\psi_{h}^{sys}$, $\overline{\psi_{h}}$, and $PCF_{h}$ from arrival time distributions. Using the same approach described in Section 3.6, we identified destination zones in Brisbane that exhibit high arrival time peakedness and compared them with the previously analysed origin zones. The spatial distribution of these destination zones is presented in Figure 19. Similar to the origin-based analysis, the results reveal notable discrepancies between destination zones with high system-wide arrival time peakedness ($\psi_{h}^{sys}$) and those with high user-level arrival time peakedness ($\overline{\psi_{h}}$). Even though destination zones in Brisbane during the AM period are generally concentrated in the city centre, there are still considerable differences between destinations with high trip concentration based on aggregated arrival distributions (first row) and those where travellers individually exhibit a high degree of peakedness in their respective arrival time distributions (second row). This reinforces our earlier finding that system-wide demand patterns do not always align with individual users’ travel time preferences. The distinct zone groups presented in the final row of Figure 19 further highlights the importance of considering both user-level peakedness ($\overline{\psi_{h}}$) and peak window coincidence ($PCF_{h}$) separately. By distinguishing between these measures, transit operators can better address the impact of destination disruptions in different scenarios, such as disruptions that impose high rescheduling or late arrival penalties on users with rigid arrival schedules (due to high $\overline{\psi_{h}}$) and those that affect a large number of travellers simultaneously by disrupting a common peak period (due to high $PCF_{h}$).
This analysis also provides additional insights into areas that exhibit high peakedness in both departure and arrival patterns by jointly analysing both origin and destination zones. These zones represent locations where rigid travel behaviour is observed both as origins and destinations. A disruption in these zones could impact a large number of people’s departures as well as a large number of people’s arrivals, amplifying the overall system-wide effect.
Figure 19: The spatial distribution of origin zones with high departure time peakedness and destination zones with high arrival time peakedness (AM period, Brisbane).